# Towards chemical accuracy for chemi- and physisorption with an efficient density functional

### v1

@@ Abstract
I Introduction
Molecular adsorption on solid surfaces is a challenging yet critical topic in surface science as a fundamental reaction step for surface chemical reactions. Understanding this processes of molecules adsorbing to solid surfaces is an essential step in understanding important surface phenomena including semiconductor processing, corrosion, electrochemistry, and heterogeneous catalysis[1]. For example, understanding how strongly a molecule binds to a catalytic surface is essential for designing improved heterogeneous catalysts, which have been estimated to drive over 20% of industrial chemistry [2] and are essential in driving the transition to carbon neutral operation[3, 4].
@@ I Introduction
Understanding molecular adsorption on solid surfaces is limited in practice by the difficulty available first-principles methods have in describing the interactions between a finite molecule and a semi-infinite solid surface[5]. This process can involve the formation of both short-range chemical bonds, termed chemisorption, and long-range physical interactions, e.g., the van der Waals (vdW) interaction, termed physisorption. Calculations using high-level methods such as quantum Monte Carlo (QMC)[6], coupled-cluster singles and doubles with perturbative triples CCSD(T)[7], and random-phase approximation (RPA)[8, 9] can capture both chemisorption and physisorption accurately but their severe computational cost prevents their routine application to molecular adsorption problems. Calculations using density functional theory (DFT) [10, 11] are comparatively cheap however, this combination of computational efficiency and useful accuracy has resulting in their becoming a widely accepted workhorse method for surface calculations. The past decades in particular have seen extraordinary progress in DFT driven computational heterogeneous catalysis [12, 13], resulting in greatly accelerated computer based catalyst design [14, 15, 16]. However, despite this wide adoption, the accuracy of efficient exchange-correlation density functional approximations (DFAs) has not yet reached a point where both chemisorption and physisorption can be reliably predicted simultaneously.
Figure 1: (a) Bivariate plot of Root Mean Square Deviation (RMSD) of the chemisorption and physisorption energies per adsorbate for CE39 systems. The blue dashed line represents the 0.25 eV RMSD threshold for both physisorption and chemisorption, with only Opt(MS+rVV10) edging this boundary. (b) Binding energy curves for graphene adsorbed in Ni(111) surface from RPBE+D3[17], BEEF-vdW [18, 19], and Opt(MS+rVV10) compared with RPA[20] results. The blue shaded regions correspond to the experimental adsorption distance [21] of $2.11\pm 0.07$ and to the interval of experimental adsorption energy [22, 23] that ranges from $74\pm 8.1$ meV.
@@
The accuracy of a DFT calculation is largely determined by the chosen exchange-correlation approximation, very many of which have been proposed. Approximate exchange-correlation functionals can broadly be categorized into a hierarchy of increasing sophistication accompanied by increasing computational cost and (ideally) increasing accuracy [24]. The generalized gradient approximation (GGA) and meta-GGA class of exchange-correlation functionals are well suited for computing surface quantities in chemistry and condensed matter physics since these approximations require only semilocal ingredients and hence remain computationally inexpensive [25, 26]. The absence of non-local ingredients means that long-range van der Waals interactions are absent from GGAs and meta-GGAs however, which can add significant errors to surface calculations [27, 28, 29, 30]. To remedy this, various vdW corrections have been developed to pair with the semilocal approximations [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. One such case, highly relevant to this work, involved pairing the meta-GGA MS2[44, 45] with the rVV10[46] vdW correction without optimizing its parameters. This resulted in a functional that failed to accurately capture the vdW wells, particularly for the well-studied H2 + Cu(111) system[47]. Typically, the vdW correction is fitted independently while the GGA or meta-GGA functional retains its parameterization based on general chemistry applications. This approach often leads to a trade-off: most GGA/meta-GGA functionals with vdW corrections perform well for either chemisorption or physisorption, but not both simultaneously. Because the semilocal functional and vdW correction are not designed in conjunction, existing GGA and meta-GGA functionals, even with vdW corrections, struggle to capture both types of adsorption with high accuracy at the same time.
@@
In this work we propose a novel density functional with balanced performance for both chemi- and physisorption by concurrently optimizing both semilocal component and the long-range vdW correction against the prototypical CO/Pt(111) and the Ar2 binding energy curves, see Section II.3 for further details. The resulting density functional, termed “Opt(MS+rVV10)” which is the topic of this work is a reparameterization of both the meta-GGA Made Simple (MS) exchange-correlation functional [44] and the rVV10 long range vdW correction [46]. Opt(MS+rVV10) shows improved and balanced performance for both the physisorption and chemisorption of molecules adsorbed on transition metal surfaces compared to the other DFAs popular for surface science [48]. Furthermore, Opt(MS+rVV10) shows the most balanced performance for molecular adsorptions of the DFAs considered and yields both the chemi- and physisorption local minima for graphene adsorbed on a Ni(111) surface, predicting a binding energy curve in close agreement with that from the high-level RPA [20].
II Results
II.1 Bivariate Plot
@@
The CE39 dataset, recently proposed by Wellendorf et al. [48], comprises experimental data for 39 well-defined adsorption energies of various gas molecules adsorbed on eight transition metal surfaces. It includes two subsets: strongly bound chemisorbed systems and weakly bound physisorbed systems where van der Waals (vdW) interactions dominate. Figure 1(a) presents a bivariate plot contrasting the performance of Opt(MS+rVV10) against other widely used density functionals, such as the local density approximation (LDA[49]), generalized gradient approximations (PBE[50] and RPBE[17]), meta-GGAs (SCAN[51], MS2[52], and M06L), vdW-corrected GGAs (RPBE+D3, BEEF-vdW[19], and optPBE-vdW), and vdW-corrected meta-GGAs (SCAN+rVV10). These functionals were selected due to their popularity in surface chemistry (e.g., RPBE) and broader applications (e.g., PBE).
@@
The LDA functional significantly overestimates chemisorption with an root mean square deviation (RMSD) around 1.0 eV but predicts physisorption with reasonable accuracy (0.4 eV RMSD). This unexpected accuracy in physisorption likely results from error cancellation, as LDA generally overestimates electron density overlap bonding while failing to account for long-range vdW interactions.PBE, a GGA functional, improves upon LDA for chemisorption, reducing the RMSD from 1.0 eV to 0.38 eV, while RPBE further lowers it to 0.22 eV. However, both functionals struggle with physisorption, exhibiting RMSDs of 0.46 eV and 0.71 eV for PBE and RPBE, respectively.
@@
When vdW corrections are applied, RPBE+D3 [53] significantly enhances RPBE’s performance, delivering a more balanced description of both chemisorption (0.31 eV RMSD) and physisorption (0.35 eV RMSD). The BEEF-vdW functional, developed using a Bayesian approach to fit exchange-correlation parameters for surface chemistry [19], achieves an even better RMSD for chemisorption (0.21 eV) compared to RPBE+D3 and a slightly improved physisorption RMSD (0.34 eV). However, this imbalance may be attributed to the fact that 17 chemisorbed systems from the CE39 dataset were included in the BEEF-vdW training set. While optPBE-vdW performs best for physisorbed systems among the tested functionals, its overestimation of chemisorption remains unsatisfactory [54].
@@
At the meta-GGA level, although SCAN [51] has been highly successful in addressing longstanding issues in condensed matter physics, such as strongly correlated cuprates and phase transitions [55, 56, 26], it overestimates chemisorption with an RMSD of approximately 0.44 eV. This may be due to self-interaction errors, which can lead to an overestimation of charge transfer between molecules and metal surfaces, as evidenced in CO/Pt(111) adsorption studies [57]. SCAN, however, has demonstrated an ability to capture intermediate vdW interactions [58], making it reasonably accurate for physisorption with an RMSD of around 0.30 eV. Another meta-GGA, M06L, performs exceptionally well for chemisorption, with an RMSD of 0.21 eV, but it is significantly less accurate for physisorption (RMSD of 0.37 eV). This is surprising, given that M06L was the first semilocal functional to successfully capture intermediate vdW interactions [59]. The discrepancy may be due to the highly empirical nature of M06L, which over-represents molecular systems, leading to an exaggerated emphasis on chemical bonding at the expense of accurately modeling physisorption.
@@
Even without a vdW correction, the MS2 meta-GGA remains one of the most balanced functionals, offering reasonable accuracy across both chemisorption and physisorption. This balance motivated the pairing of MS2 with the rVV10 vdW correction, alongside a re-optimization of the internal parameters, to enhance accuracy while maintaining balanced performance. As demonstrated, the resulting Opt(MS+rVV10) functional provides the best balanced performance for both chemisorption (RMSDsimilar-$\sim$ 0.24 eV) and physisorption (RMSDsimilar-$\sim$ 0.26 eV).
II.2 Graphene adsorption on Ni(111)
@@
Graphene adsorption on metal is highly metal dependent forming strong chemical bonding with some metals and only weak van der Waals interaction with others. Graphene on Ni has shown exceptional electronic properties in semiconducting technology [60], however understanding the interactions between graphene and Ni(111) surface has been challenging for both experiment and theory. A study using angle resolved photoemission (ARPES) reveals strong chemical bonding combined with weak vdW interactions between the graphene sheet and Ni(111) [61]. Several theoretical studies have been made drawing conflicting conclusions; some predicting dominant chemisorption minima and others predicting deeper physisorption minima[62, 63, 64, 65, 66]. High level RPA calculations have established that both chemi- and phsysisorption minima should have similar depth, in good accord with the experimentally determined binding energy, as shown in Figure 1(b). Opt(MS+rVV10) predicts chemisorption and physisorption minima at 2.15 Å and 3.5 Å, respectively, compared to RPA@PBE values of 2.17 Å and 3.27 Å. The binding energies for these minima are -58 meV and -57 meV for Opt(MS+rVV10), versus -70 meV and -67 meV for RPA@PBE.
@@
The binding energy curves of graphene adsorbed on Ni surface calculated using: RPBE+D3, BEEF-vdW, and Opt(MS+rVV10), are shown in Figure 1(b). These functionals were chosen as representative vdW corrected functionals that exhibit good balance between chemi- and physisorption in Figure 1(a). The binding curves for BEEF-vdW and RPA@PBE are taken from Refs 18 and 20, respectively. Low-energy electron diffraction (LEED)[21] experiments measure the equilibrium separation of the graphene sheet on Ni(111) as $2.11\pm 0.07$, which is shown in Figure 1(b) as vertical dashed blue line with highlighted uncertainty. The corresponding binding energy has been measured by auger spectroscopy [22, 23] as $74\pm 8.1$ meV, illustrated by the horizontal dashed blue line with highlighted uncertainty.
Figure 2: (a) Heat map of 3rd order polynomial function fitted to Absolute Errors of CO adsorbed on Pt(111) surface as a function of $\kappa$ and $b$ fitting parameters, Errors are in eV. The minimizing parameters are marked in red as $\kappa=0.2501$ and $b=0.9104$. (b) Heat map of normalized absolute errors of atomization energies subset AE6[67]
@@
Figure 1(b) shows that RPA@PBE predicts a chemisorption minimum with both the equilibrium position and the binding energy in good agreement with experimental data. Several studies and RPA@PBE support that graphene on Ni(111) has an additional minimum due to strong chemical and physical adsorption present in the graphene on Ni(111) system [20, 22, 23]. BEEF-vdW accurately predicts chemisorption separation distance, though the binding energy is greatly underestimated. Surprisingly, BEEF-vdW predicts a long-range physisorption minimum that is much deeper than its chemisorption minimum. Furthermore, RPBE+D3 shows a clear physisorption minimum at the expected separation, but the short-range chemisorption minimum is absent. Finally, the newly optimized Opt(MS+rVV10) functional agrees well by having a double minimum similar to RPA@PBE. Even though the binding energy predicted by the Opt(MS+rVV10) is slightly higher than the RPA@PBE and experimental uncertainties, the minimum intermolecular distance of chemisorption and physisorption along with the binding energies show the closest agreement to the benchmark and experimental data.
II.3 Determination of parameters for Opt(MS+rVV10)

### v2

# {tocentry}[Uncaptioned image]

@@ Abstract
Key Words: Adsorption energy, Heterogeneous catalysis, Density functional theory, meta-generalized gradient approximation (meta-GGA), van der Waals interaction
1 Introduction
Molecular adsorption on solid surfaces is a challenging yet critical topic in surface science as a fundamental reaction step for surface chemical reactions. Understanding this processes of molecules adsorbing to solid surfaces is an essential step in understanding important surface phenomena including semiconductor processing, corrosion, electrochemistry, and heterogeneous catalysis1. In catalysis, adsorption energies are particularly significant because they directly influence reaction rates, selectivity, and the stability of intermediates, all of which are essential for efficient catalyst design. Predicting adsorption energies accurately is vital for optimizing catalytic activity, as they determine stable intermediates, accessible transition states, and, ultimately, guide the design of active, selective, and durable catalytic materials. Misestimations in adsorption energy can lead to inaccurate predictions, hindering the development of improved catalysts essential for industrial and sustainable applications.
@@ 1 Introduction
However, accurately understanding molecular adsorption on solid surfaces is challenging due to the limitations of available first-principles methods in describing interactions between finite molecules and semi-infinite solid surfaces2. This process can involve the formation of both short-range chemical bonds, termed chemisorption, and long-range physical interactions, e.g., the van der Waals (vdW) interaction, termed physisorption. Calculations using high-level methods such as quantum Monte Carlo (QMC)3, coupled-cluster singles and doubles with perturbative triples CCSD(T)4, and random-phase approximation (RPA)5, 6 can capture both chemisorption and physisorption accurately but their severe computational cost prevents their routine application to molecular adsorption problems. Calculations using density functional theory (DFT) 7, 8 are comparatively cheap however, this combination of computational efficiency and useful accuracy has result ing in their becoming a widely accepted workhorse method for surface calculations. The past decades in particular have seen extraordinary progress in DFT driven computational heterogeneous catalysis 9, 10, resulting in greatly accelerated computer based catalyst design 11, 12, 13. Nevertheless, the limitations of exchange-correlation density functional approximations (DFAs) prevent DFT from reliably predicting both chemisorption and physisorption interactions with equal accuracy, thus posing a significant challenge in catalytic material design.
@@
The accuracy of a DFT calculation is largely determined by the chosen exchange-correlation approximation, very many of which have been proposed. Approximate exchange-correlation functionals can broadly be categorized into a hierarchy of increasing sophistication accompanied by increasing computational cost and (ideally) increasing accuracy 14. The generalized gradient approximation (GGA) and meta-GGA class of exchange-correlation functionals are well suited for computing surface quantities in chemistry and condensed matter physics since these approximations require only semilocal ingredients and hence remain computationally inexpensive 15, 16. The absence of non-local ingredients means that long-range van der Waals interactions are absent from GGAs and meta-GGAs however, which can add significant errors to surface calculations 17, 18, 19, 20. To remedy this, various vdW corrections have been developed to pair with the semilocal approximations 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33. One such case, highly relevant to this work, involved pairing the meta-GGA MS234, 35 with the rVV1036 vdW correction without optimizing its parameters. This resulted in a functional that failed to accurately capture the vdW wells, particularly for the well-studied H2 + Cu(111) system37. Typically, the vdW correction is fitted independently while the GGA or meta-GGA functional retains its parameterization based on general chemistry applications. This approach often leads to a trade-off: most GGA/meta-GGA functionals with vdW corrections perform well for either chemisorption or physisorption, but not both simultaneously. Because the semilocal functional and vdW correction are not designed in conjunction, existing GGA and meta-GGA functionals, even with vdW corrections, struggle to capture both types of adsorption with high accuracy at the same time. Recent advances in many-body dispersion methods, particularly the work by Tkatchenko and colleagues on screened van der Waals interactions and many-body dispersion (MBD) effects32, 38, 39, have demonstrated the importance of collective electronic effects in surface adsorption beyond pairwise additivity.
@@
In this work we propose a novel density functional with balanced performance for both chemi- and physisorption by concurrently optimizing both semilocal component and the long-range vdW correction against the prototypical CO/Pt(111) and the Ar2 binding energy curves, see Section 2.3 for further details. The resulting density functional, termed “Opt(MS+rVV10)” which is the topic of this work is a reparameterization of both the meta-GGA Made Simple (MS) exchange-correlation functional 34 and the rVV10 long range vdW correction 36. Opt(MS+rVV10) shows improved and balanced performance for both the physisorption and chemisorption of molecules adsorbed on transition metal surfaces compared to the other DFAs popular for surface science 40. Furthermore, Opt(MS+rVV10) shows the most balanced performance for molecular adsorptions of the DFAs considered and yields both the chemi- and physisorption local minima for graphene adsorbed on a Ni(111) surface, predicting a binding energy curve in close agreement with that from the high-level RPA 41.
2 Results
2.1 Bivariate Plot
@@
The CE39 dataset, proposed by Wellendorf et al. 40, comprises experimental data for 39 well-defined adsorption energies of various gas molecules adsorbed on eight transition metal surfaces. These experimental adsorption energies have been corrected for zero-point vibrational energy (ZPVE) and RT contributions, following the procedure outlined in Ref.40, to approximate 0 K enthalpies. This correction enables a consistent and meaningful comparison to static-lattice DFT energies, which are calculated at 0 K without thermal contributions. It includes two subsets: strongly bound chemisorbed systems and weakly bound physisorbed systems where van der Waals (vdW) interactions dominate. Figure 1(a) presents a bivariate plot contrasting the performance of Opt(MS+rVV10) against other widely used density functionals, such as the local density approximation (LDA42), generalized gradient approximations (PBE43 and RPBE44), meta-GGAs (SCAN45, MS246, and M06L), vdW-corrected GGAs (RPBE+D3, BEEF-vdW47, and optPBE-vdW), and vdW-corrected meta-GGAs (SCAN+rVV10). These functionals were selected due to their popularity in surface chemistry (e.g., RPBE) and broader applications (e.g., PBE).
Figure 1: (a) Bivariate plot of Root Mean Square Deviation (RMSD) of the chemisorption and physisorption energies per adsorbate for CE39 systems. The blue dashed line represents the 0.25 eV RMSD threshold for both physisorption and chemisorption, with only Opt(MS+rVV10) edging this boundary. (b) Binding energy curves for graphene adsorbed in Ni(111) surface from RPBE+D344, BEEF-vdW 48, 47, and Opt(MS+rVV10) compared with RPA41 results. The blue dotted lines indicate the experimental adsorption distance 49 of $2.11\pm 0.07$ Å, representing the uncertainty range in the measurement, while the orange shaded regions correspond to the experimental adsorption energy range 50, 51 of $74\pm 8.1$ meV, reflecting the uncertainty in the reported values.
@@
The LDA functional significantly overestimates chemisorption with an root mean square deviation (RMSD) around 1.0 eV but predicts physisorption with reasonable accuracy (0.4 eV RMSD). This unexpected accuracy in physisorption likely results from error cancellation, as LDA generally overestimates electron density overlap bonding while failing to account for long-range vdW interactions.PBE, a GGA functional, improves upon LDA for chemisorption, reducing the RMSD from 1.0 eV to 0.38 eV, while RPBE further lowers it to 0.22 eV. However, both functionals struggle with physisorption, exhibiting RMSDs of 0.46 eV and 0.71 eV for PBE and RPBE, respectively. More recently, Sharada et al.52 extended the CE39 set by adding two additional adsorption reactions, creating the AD41 benchmark, which has been used to assess new density functional approximations such as the empirically fitted meta-GGA MCML53. This functional, while grounded in physical constraints and informed by experimental and quantum chemistry reference data, is empirically optimized to yield improved accuracy for surface and gas-phase reaction energetics.
@@
When vdW corrections are applied, RPBE+D3 54 significantly enhances RPBE’s performance, delivering a more balanced description of both chemisorption (0.31 eV RMSD) and physisorption (0.35 eV RMSD). The BEEF-vdW functional, developed using a Bayesian approach to fit exchange-correlation parameters for surface chemistry 47, achieves an even better RMSD for chemisorption (0.21 eV) compared to RPBE+D3 and a slightly improved physisorption RMSD (0.34 eV). However, this imbalance may be attributed to the fact that 17 chemisorbed systems from the CE39 dataset were included in the BEEF-vdW training set. While optPBE-vdW performs best for physisorbed systems among the tested functionals, its overestimation of chemisorption remains unsatisfactory 55.
@@
At the meta-GGA level, although SCAN 45 has been highly successful in addressing longstanding issues in condensed matter physics, such as strongly correlated cuprates and phase transitions 56, 57, 16, it overestimates chemisorption with an RMSD of approximately 0.44 eV. This may be due to self-interaction errors, which can lead to an overestimation of charge transfer between molecules and metal surfaces, as evidenced in CO/Pt(111) adsorption studies 58. SCAN, however, has demonstrated an ability to capture intermediate vdW interactions 59, making it reasonably accurate for physisorption with an RMSD of around 0.30 eV. Another meta-GGA, M06L, performs exceptionally well for chemisorption, with an RMSD of 0.21 eV, but it is significantly less accurate for physisorption (RMSD of 0.37 eV). This is surprising, given that M06L was the first semilocal functional to successfully capture intermediate vdW interactions 60. The discrepancy may be due to the highly empirical nature of M06L, which over-represents molecular systems, leading to an exaggerated emphasis on chemical bonding at the expense of accurately modeling physisorption.
@@
[…]
2.2 Graphene adsorption on Ni(111)
@@
Graphene adsorption on metal is highly metal dependent forming strong chemical bonding with some metals and only weak van der Waals interaction with others. Graphene on Ni has shown exceptional electronic properties in semiconducting technology 61, however understanding the interactions between graphene and Ni(111) surface has been challenging for both experiment and theory. A study using angle resolved photoemission (ARPES) reveals strong chemical bonding combined with weak vdW interactions between the graphene sheet and Ni(111) 62. Several theoretical studies have been made drawing conflicting conclusions; some predicting dominant chemisorption minima and others predicting deeper physisorption minima63, 64, 65, 66, 67. High level RPA calculations have established that both chemi- and phsysisorption minima should have similar depth, in good accord with the experimentally determined binding energy, as shown in Figure 1(b). Opt(MS+rVV10) predicts chemisorption and physisorption minima at 2.15 Å and 3.5 Å, respectively, compared to RPA@PBE values of 2.17 Å and 3.27 Å. The binding energies for these minima are -58 meV and -57 meV for Opt(MS+rVV10), versus -70 meV and -67 meV for RPA@PBE.
@@
The binding energy curves of graphene adsorbed on Ni surface calculated using: RPBE+D3, BEEF-vdW, and Opt(MS+rVV10), are shown in Figure 1(b). These functionals were chosen as representative vdW corrected functionals that exhibit good balance between chemi- and physisorption in Figure 1(a). The binding curves for BEEF-vdW and RPA@PBE are taken from Refs 48 and 41, respectively. Low-energy electron diffraction (LEED)49 experiments measure the equilibrium separation of the graphene sheet on Ni(111) as $2.11\pm 0.07$ Å, which is shown in Figure 1(b) as vertical dashed blue line with highlighted uncertainty. The corresponding binding energy has been measured by auger spectroscopy 50, 51 as $74\pm 8.1$ meV, illustrated by the horizontal dashed blue line with highlighted uncertainty.
@@
Figure 1(b) shows that RPA@PBE predicts a chemisorption minimum with both the equilibrium position and the binding energy in good agreement with experimental data. Several studies and RPA@PBE support that graphene on Ni(111) has an additional minimum due to strong chemical and physical adsorption present in the graphene on Ni(111) system 41, 50, 51. BEEF-vdW accurately predicts chemisorption separation distance, though the binding energy is greatly underestimated. Surprisingly, BEEF-vdW predicts a long-range physisorption minimum that is much deeper than its chemisorption minimum. Furthermore, RPBE+D3 shows a clear physisorption minimum at the expected separation, but the short-range chemisorption minimum is absent. Finally, the newly optimized Opt(MS+rVV10) functional agrees well by having a double minimum similar to RPA@PBE. Even though the binding energy predicted by the Opt(MS+rVV10) is slightly higher than the RPA@PBE and experimental uncertainties, the minimum intermolecular distance of chemisorption and physisorption along with the binding energies show the closest agreement to the benchmark and experimental data. Even though the binding energy predicted by the Opt(MS+rVV10) is slightly higher than the RPA@PBE and experimental uncertainties, the minimum intermolecular distance of chemisorption and physisorption along with the binding energies show the closest agreement to the benchmark and experimental data. While our approach does not explicitly include many-body dispersion effects as implemented in advanced methods 68, 69, the concurrent optimization of both the semilocal MS functional and rVV10 parameters may implicitly capture some collective effects. The MS2 meta-GGA’s known ability to partially capture intermediate-range vdW interactions 46, combined with surface-specific parameterization, allows for some incorporation of the cooperative electronic response that characterizes many-body effects. In terms of computational efficiency, RPBE+D3 is the fastest among the tested vdW-corrected functionals. Opt(MS+rVV10) is approximately twice as expensive as RPBE+D3, whereas BEEF-vdW and SCAN+rVV10 are about four times more expensive. This trade-off between accuracy and computational cost may be an important consideration for large-scale or high-throughput studies.
Figure 2: (a) Heat map of 3rd order polynomial function fitted to Absolute Errors of CO adsorbed on Pt(111) surface as a function of $\kappa$ and $b$ fitting parameters, Errors are in eV. The minimizing parameters are marked in red as $\kappa=0.2501$ and $b=0.9104$. (b) Heat map of normalized absolute errors of atomization energies subset AE670
2.3 Determination of parameters for Opt(MS+rVV10)